Solve for 'a'

Determine the values of a for which the following system of linear equations has no solutions, a unique solution, or infinitely many solutions. You can select 'always', 'never', 'a = ', or 'a ≠', then specify a value or comma-separated list of values.


2x1−4x2+2x3 = 6
−3x1+6x2−2x3 = 11
ax1−6x2+6x3 = 4


So far I was able to change it into matrix 


2     -4     2    |   6

-3     6     -2   |   -11

a      -6    6    |   4


What would be my following sets to solve for 'a'?
Answer #1

Rgt Rg + 16+20) Ra 2a,- udat ada=6 -321+622- 282=-11 an, - 682 + 642=4 Ax=b 6 lo 4 ३ - Il . A = -3 6 - a - G 6 its Augmentedil Rank (A) = Rank (C) then = no of unknowns then system has then system has unique solh for Rand unique solh. 89.90 90 ie 39